If the area of a square is 68cm, find the length of one side? Show all work.
Answer by stanbon(75887)

If the area of a square is 68cm, find the length of one side? Show all work.
-------------------
Area of a square = (side)^2
-----
Equation:
68 = s^2
----
s = sqrt(68) = 8.2462 cm
=====================
Cheers,
Stan H.
=====================